# NAI2: Learning Noise-Aware Illumination Interpolator for Unsupervised Low-Light Image Enhancement

### v3

@@ 5.2 Low-light Image Enhancement
Benchmark Datasets and Evaluation Metrics. To rigorously evaluate the effectiveness of our proposed method in low-light image enhancement, we conducted assessments across multiple datasets. The evaluation encompassed well-established datasets such as MIT [29] and LOL [15], renowned for their diverse light condition images. The MIT dataset contains noiseless low-light images, serving as a benchmark for comparison, while the LOL dataset includes images captured under various light conditions, often exhibiting significant noise. Additionally, we included several unpaired datasets including DICM [30], LIME [31], MEF [32], NPE [33] and VV [34] for testing directly. In evaluating the efficacy of our method, we employed a comprehensive set of evaluation metrics. This includes two full-reference metrics, PSNR and SSIM [35], known for their precision in assessing image quality. Furthermore, we incorporated no-reference metrics, DE [36], EME [37], LOE [38] and NIQE [39], essential for evaluating images in the absence of reference images.
@@
Benchmark Evaluation. We performed quantitative and qualitative evaluations of the proposed method against several state-of-the-art LLIE methods, including traditional methods (i.e. LIME [31] and Structure-Revealing Retinex Model [40] ), supervised learning methods (i.e., Retinex [15], GLADNet [17], DRBN [18] and MBLLEN [16]), and unsupervised learning methods (i.e, EnGAN [9], ZeroDCE [5], RetinexDIP [10], RUAS [7], and SCI  [8] ). Presented below are the experimental findings following the outlined experimental conditions and settings.
Qualitative evaluation. Qualitative results are displayed in Figure 9. As illustrated, previous efforts did not achieve the desired enhancement results, resulting in inconspicuous details, unnatural colors and overexposure or underexposure. In contrast, our NAI2 achieves the best visual quality with vivid colors and outstanding texture details. In contrast, our approach is both fast and flexible, and delivers highly effective results. We achieve real-time enhancement without the need for external hyperparameters finetuning or early stop strategies, and demonstrate superior performance compared to existing methods.
@@ 5.3 Ablation Studies
Effects of SCD and LII. Table 5 presents the results of network module (i.e., SCD and LII) ablation on the LOL dataset. It should be noted that in the table, [M1] refers to the model without the SCD module, and [M2] represents the removal of the LII module, which is replaced by the unsupervised illumination learning network proposed in the recent work SCI [8]. We observe that without the SCD module, the performance decreased by $2.1901dB$ and $0.2008$ on the PSNR and SSIM metrics, respectively. Moreover, the substitution of LII for SCI significantly attenuates the performance on all metrics, especially with a $5.1061dB$ decrease in PSNR. We argue that the root cause of this phenomenon is that the substituted Model [M2] does not comply with the loss optimization procedure under the illumination prior constraint condition with a simple loss function. In the other hand, it indicates that our LII module satisfy the illumination prior naturally.
@@
Exploration on Reference-Free Loss. Our study investigates the effectiveness of self-regularized recovery loss in comparison with popular unsupervised learning loss functions, including $\mathcal{L}_{DCE}$ from ZeroDCE [5] and $\mathcal{L}_{SCI}$ from SCI [8], as shown in Table 6. We built Model [M5] and [M6] separately using the corresponding complex loss configurations. Hyperparameters of the loss functions are utilized without any coefficient adjustment for evaluate its robustness. However, due to the lack of a denoising function of these model, we remove the $\mathcal{L}_{nr}$ as well as SCD term and construct Model [M7] for justice. None of the models include an SCD module. Our quantitative results show a substantial improvement in Model [M7] over Models [M5] and [M6], such as a $67.6\%$ PSNR score improvement, demonstrating robustness and effectiveness in color restoration, as indicated in Figure 11 and Table 6. Additionally, we explore self-adaptable noise removal loss by removing the adaptive factor $\sigma(\mathbf{x})$ from $\mathcal{L}_{nr}$ to train Model [M8], which leads to the degradation of the loss to the most common denoise loss $\mathcal{L}_{TV}$ [27]. Our final model shows improvements in the PSNR and SSIM metrics of $0.2111dB$ and $0.1160$, respectively.

### v4

@@ 5.2 Low-light Image Enhancement
Benchmark Datasets and Evaluation Metrics. To rigorously evaluate the effectiveness of our proposed method in low-light image enhancement, we conducted assessments across multiple datasets. The evaluation encompassed well-established datasets such as MIT bychkovsky2011learning and LOL wei2018deep , renowned for their diverse light condition images. The MIT dataset contains noiseless low-light images, serving as a benchmark for comparison, while the LOL dataset includes images captured under various light conditions, often exhibiting significant noise. Additionally, we included several unpaired datasets including DICM DICM , LIME LIME , MEF MEF , NPE NPE and VV VV for testing directly. In evaluating the efficacy of our method, we employed a comprehensive set of evaluation metrics. This includes two full-reference metrics, PSNR and SSIM wang2004image , known for their precision in assessing image quality. Furthermore, we incorporated no-reference metrics, DE DE , EME EME , LOE LOE and NIQE NIQE , essential for evaluating images in the absence of reference images.
@@
Benchmark Evaluation. We performed quantitative and qualitative evaluations of the proposed method against several state-of-the-art LLIE methods, including traditional methods (i.e. LIME LIME and Structure-Revealing Retinex Model SRRM ), supervised learning methods (i.e., Retinex wei2018deep , GLADNet wang2018gladnet , DRBN yang2020fidelity and MBLLEN lv2018mbllen ), and unsupervised learning methods (i.e, EnGAN jiang2021enlightengan , ZeroDCE guo2020zero , RetinexDIP zhao2021retinexdip , RUAS liu2021retinex , and SCI  ma2022toward ). Presented below are the experimental findings following the outlined experimental conditions and settings.
Qualitative evaluation. Qualitative results are displayed in Figure 9. As illustrated, previous efforts did not achieve the desired enhancement results, resulting in inconspicuous details, unnatural colors and overexposure or underexposure. In contrast, proposed method achieves the best visual quality with vivid colors and outstanding texture details. In contrast, our approach is both fast and flexible, and delivers highly effective results. We achieve real-time enhancement without the need for external hyperparameters finetuning or early stop strategies, and demonstrate superior performance compared to existing methods.
@@ 5.3 Ablation Studies
Effects of SCD and LII. Table 5 presents the results of network module (i.e., SCD and LII) ablation on the LOL dataset. It should be noted that in the table, [M1] refers to the model without the SCD module, and [M2] represents the removal of the LII module, which is replaced by the unsupervised illumination learning network proposed in the recent work SCI ma2022toward . We observe that without the SCD module, the performance decreased by $2.1901dB$ and $0.2008$ on the PSNR and SSIM metrics, respectively. Moreover, the substitution of LII for SCI significantly attenuates the performance on all metrics, especially with a $5.1061dB$ decrease in PSNR. We argue that the root cause of this phenomenon is that the substituted Model [M2] does not comply with the loss optimization procedure under the illumination prior constraint condition with a simple loss function. In the other hand, it indicates that our LII module satisfy the illumination prior naturally.
@@
Exploration on Reference-Free Loss. Our study investigates the effectiveness of self-regularized recovery loss in comparison with popular unsupervised learning loss functions, including $\mathcal{L}_{DCE}$ from ZeroDCE guo2020zero and $\mathcal{L}_{SCI}$ from SCI ma2022toward , as shown in Table 6. We built Model [M5] and [M6] separately using the corresponding complex loss configurations. Hyperparameters of the loss functions are utilized without any coefficient adjustment for evaluate its robustness. However, due to the lack of a denoising function of these model, we remove the $\mathcal{L}_{nr}$ as well as SCD term and construct Model [M7] for justice. None of the models include an SCD module. Our quantitative results show a substantial improvement in Model [M7] over Models [M5] and [M6], such as a $67.6\%$ PSNR score improvement, demonstrating robustness and effectiveness in color restoration, as indicated in Figure 11 and Table 6. Additionally, we explore self-adaptable noise removal loss by removing the adaptive factor $\sigma(\mathbf{x})$ from $\mathcal{L}_{nr}$ to train Model [M8], which leads to the degradation of the loss to the most common denoise loss $\mathcal{L}_{TV}$ rudin1992nonlinear . Our final model shows improvements in the PSNR and SSIM metrics of $0.2111dB$ and $0.1160$, respectively.